Show that z = i is a root of the equation z
+ z + z − 1 = 0. Find the three
other roots.
1
Expert's answer
2020-05-25T21:11:57-0400

Putting "z=i" in the following term "z^{4}+z^{3}+z-1" , we have

"i^{4}+i^{3}+i-1=1-i+i-1=0"

thus "z=i" is a root of the equation "z^{4}+z^{3}+z-1=0"

Since the equation "z^{4}+z^{3}+z-1=0" has the root "z=i" , then by using the division rule, we have that

"z^{4}+z^{3}+z-1=(z-i)(z^{3}+(i+1)z^{2}+(i-1)z-i)"

Since "z=i" is a root of the equation "z^{4}+z^{3}+z-1=0" , then "z=-i" is also a root, thus


"z^{3}+(i+1)z^{2}+(i-1)z-i=(z+i)(z^{2}+z-1)"

the equation "z^{2}+z-1=0" can be solved by the general law of the quadratic function , the roots of this equation is are "z= (-1+\sqrt5 )/2" and "z= (-1-\sqrt5 )/2"


the four roots are

"z=i" , "z=-i" , "z= (-1+\sqrt5 )/2" and "z= (-1-\sqrt5 )/2"